A sample of size 47 will be drawn from a population with mean 28 and standard deviation 6. Find the probability that x will be greater than 29.
Answer #1

solution:-

given that mean = 28 , standard deviation = 6

and sample n = 47

formula

z = (x-mean)/(standard deviation/sqrt(n))

here asked that

=>P(x > 29)

= P(z > (29-28)/(6/sqrt(47)))

= P(z > 1.14)

= 1 - P(z < 1.14)

= 1 - 0.8729

= 0.1271

note: z value from standard normal table